OpenStudy (anonymous):

For the equation f(x)=arcsinx find the Taylor Polynomial of degree 3 of at c=1/2

OpenStudy (amistre64):

sin^-1 derivatives.... thats gonna be messy i think

OpenStudy (amistre64):

4 iterations of it eh...

OpenStudy (amistre64):

well, 3

OpenStudy (amistre64):

y = sin^1(x) (y' = 1-x^2)^(-1/2)

OpenStudy (anonymous):

f(x)=arcsin(x) f'(x)=1/Sqrt[1-x^2] f''(x)=x/(1-x^2)^(3/2)

OpenStudy (anonymous):

f(x)+f'(x)(x-c)+f''(x)((x-c)^2)/2!

OpenStudy (amistre64):

y'' = x(1-x^2)^(-3/2) yeah, that was easier than i expected :)

OpenStudy (amistre64):

y''' = 3x^2 (1-x^2)^(-5/2) right?

OpenStudy (anonymous):

so all you are doing is taking the derivative of the the equation to the 3rd prime?

OpenStudy (anonymous):

\[f(x)+f'(x)\frac{(x-c)}{1!}+f\text{''}(x)\frac{(x-c)^2}{2!!}+f\text{''}(x)\frac{(x-c)^3}{3!!}\]

OpenStudy (amistre64):

\[\sin^{-1}(x) + (1-x^2)^{-1/2} [x-(1/2)] + 2x(1-x^2)^{-3/2} [x-(1/2)]^2 + 3x^2 (1-x^2)^{-5/2}[x-(1/2)]^3\]

OpenStudy (amistre64):

i forgot the factorials..... i loathe typing these things out lol

OpenStudy (anonymous):

That's why I typed them in mathematica and copy them here

OpenStudy (anonymous):

whats the difference between your two answers?

OpenStudy (amistre64):

well; mine is missing factorials under them for starters lol

OpenStudy (anonymous):

actually same, amistreo actually plug in for f(x),f'(x)

OpenStudy (amistre64):

sin−1(x) +(1−x^2)^(−1/2) [x−(1/2)] +2x(1−x^2)^(−3/2) [x−(1/2)]^2 ---------------------------- 2! +3x^2 (1−x2)^(−5/2) [x−(1/2)]^3 ------------------------------- 3!

OpenStudy (anonymous):

Oh ok makes sense thanks so from here how do i determine the accuracy of this polynomial?

OpenStudy (amistre64):

htat i got no idea about; i just read the material on HOW to do them the other day :)

OpenStudy (anonymous):

There is actually a formula that's pain to solve, let me find it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It would be the accuracy at x=\[\sqrt{2}/2\]

OpenStudy (anonymous):

Oh, that's easy if you know x

OpenStudy (anonymous):

sorry to make you scramble

OpenStudy (anonymous):

arcsin(Sqrt[2]/2)

OpenStudy (anonymous):

sin of what angle is Sqrt[2]/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dunno its part of the same question...first it askes to find the taylor polynoial like you did than it said find the accuracy of the polynomial at x=sqrt2/2

OpenStudy (anonymous):

So sin(45degree or radian pi/4) is Sqrt(2)/2 which means arcsin(Sqrt(2)/2) is 45degree or radian pi/4.

OpenStudy (anonymous):

To find accuracy \[\left| pi/4 - what u got up there \right|\]

OpenStudy (anonymous):

oh ok gotcha

OpenStudy (anonymous):

There are other kind of question where they ask you to find error bound, which is pain to solve

OpenStudy (anonymous):

oh yea we haven't gotten that far yet.

OpenStudy (anonymous):

oh you will and you will hate it